The director of admissions at Kinzua University in Nova Scotia estimated the distribution of student admissions for the fall semester on the basis of past experience.

Admissions Probability
1,100 .6
1,400 .3
1,300 .1


Click here for the Excel Data File

  1. What is the expected number of admissions for the fall semester?

  1. Compute the variance and the standard deviation of the number of admissions. (Round your standard deviation to 2 decimal places.)
Answer #1
x P(x) xP(x) x2P(x)
1100 0.6000 660.0000 726000.00
1400 0.3000 420.0000 588000.00
1300 0.1000 130.0000 169000.00
total 1210.0000 1483000.00
E(x) =μ= ΣxP(x) = 1210.00
E(x2) = Σx2P(x) = 1483000.00
Var(x)=σ2 = E(x2)-(E(x))2= 18900.00
std deviation=         σ= √σ2 = 137.48

1) expected number of admissions for the fall semester =1210

2) variance =18900

stndard deviation =137.48